OpenStudy (anonymous):

Which of the following best defines ? (1 point) Cube root of 4 Square root of 4 Cube root of 2 Square root of 2 can someone help and tell me how to do it?

OpenStudy (dtan5457):

Defines what?

OpenStudy (anonymous):

2 2/3

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2^{2/3} }\) ?

OpenStudy (dtan5457):

More than likely is ^

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \color{red}{\rm a} ^{\color{green}{\LARGE\rm b} /\color{blue}{\LARGE\rm c} } ~~~~\Longrightarrow~~~~\sqrt[\color{blue}{\LARGE\rm c}]{ \color{red}{\rm a}^\color{green}{\LARGE\rm b}} }\)

OpenStudy (solomonzelman):

use this rule in your case.

OpenStudy (anonymous):

thats just an html code

OpenStudy (solomonzelman):

that is latex.... not that it really matters what that is.

OpenStudy (solomonzelman):

anyways, so what are you coming up with as far as the math part of this problem goes?

OpenStudy (anonymous):

is it the square root of 4?

OpenStudy (solomonzelman):

close, it is not a SQUARE root of a 4.

OpenStudy (anonymous):

Oh a cube root but how did you find that?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \color{red}{\rm a} ^{\color{green}{\LARGE\rm b} /\color{blue}{\LARGE\rm c} } ~~~~\Longrightarrow~~~~\sqrt[\color{blue}{\LARGE\rm c}]{ \color{red}{\rm a}^\color{green}{\LARGE\rm b}} }\) \(\large\color{black}{ \displaystyle \color{red}{\rm 2} ^{\color{green}{\LARGE\rm 2} /\color{blue}{\LARGE\rm 3} } ~~~~\Longrightarrow~~~~\sqrt[\color{blue}{\LARGE\rm 3}]{ \color{red}{\rm 2}^\color{green}{\LARGE\rm 2}} }\)

OpenStudy (solomonzelman):

yes, cube root of 4 is correct

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

anytime !